OpenStudy (anonymous):

Middle- earth is a fictional land inhabited by hobbits, elves, dwarves and men. The hobbits and elves are peaceful creatures that prefer slow, silent lives and appreciate nature and art. The dwarves and the men engage in physical games. The game is as follows. A tournament is one where out of the two teams that play a match, the one that loses get eliminated. The matches are played in different rounds, where in every round; half of the teams get eliminated from the tournament. If there are 8 rounds played in knock out tournament, how many matches were played?

OpenStudy (anonymous):

whats knock out tournament

OpenStudy (anonymous):

there was one match played in the 8th round between the finalists there were 2 matches played in the semifinals there were 4 matches played in the quarter finals and so on.

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

options are a) 257 b) 256 c) 72 d) 255

OpenStudy (anonymous):

Well, since there is one winner you know in round 8 only two teams remain, which means in the previous round, there were four teams (since if there were more, than more than two teams would remain), before that there must have been 8 teams for the same reason, and so on until you get back to round one. so: round 8: 2 round 7: 4 round 6: 8 round 5: 16 round 4: 32 round 3: 64 round 2: 128 round 1: 256 In numerical terms, this is your general equation. T=2^(r) where T=amount of teams in the tournament and r=amount of rounds. More specifically to your case, T=2^(8) T=256 So, your final answer is at the beginning of the tournament, there were 256 teams.

OpenStudy (anonymous):

yes got it thanx

OpenStudy (anonymous):

in round 8, 1 match was played in round 7, 2 matches were played in round 6, 4 matches were played in round 5, 8 matches were played in round 4, 16 matches were played in round 3, 32 matches were played in round 2, 64 matches were played in round 1, 128 matches were played add em all up. ignore my previous answer. I think you will get 255

OpenStudy (anonymous):

My mistake, I misread the question, Dhatraditya's answer is the correct one.

OpenStudy (anonymous):

i dont understand , if there are 4 teams in round 7, then who plays who?

OpenStudy (anonymous):

so in round 7 there are 4 teams, how do they play against each other? 2 against 2?

OpenStudy (anonymous):

suppose teams A and C were in the finals. suppose teams A,B,C,D were in semifinals, that means that A played B and won, which eliminated B and C played D and won, eliminating D

OpenStudy (anonymous):

so you have to just assume that they were matched properly

OpenStudy (anonymous):

in every round , for every match, it is 1 team against 1 team

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

and the team that loses gets eliminated from the tournament.

OpenStudy (anonymous):

so its like basketball , the college games

OpenStudy (anonymous):

, it goes finals, semi finals, quarterfinals, ...

OpenStudy (anonymous):

like this http://www.slate.com/id/2063215/

OpenStudy (anonymous):

exactly. its the same concept as a march madness bracket

OpenStudy (anonymous):

great job ;)

OpenStudy (anonymous):

thanks :)

OpenStudy (nilankshi):

good